Given g(x) = 8f(x + 2) + 12 and f(5) Fill in the blanks: Round answers to 2 decimal places as needed 90
Answer #1

Sol- f 15)=-4 g(x)= 8 f(x+2)+12 X=3 9(3)8 if (372) +12 963) = 8 $65) +12 963) = 8 (-4):+12 913) 3220answer: g(3)=–20